OpenStudy (anonymous):

Simplify the expression. 6(x - 3) - 4(x + 1) =

OpenStudy (agent47):

6x-6*3-4x-4*1. Simplify..

OpenStudy (anonymous):

Step by step: 6(x - 3) - 4 (x + 1) = 6x - 18 - 4x - 4 <----------- (distributed 6 and 4) = 6x - 4x - 18 - 4 <------------ (grouped) = 2x - 22 = 2(x-11)